OpenStudy (ihannah_banana7):

Use integration by parts, together with the techniques of this section, to evaluate the integral. (Use C for the constant of integration.)

OpenStudy (ihannah_banana7):

\[\int\limits_{}^{}9\ln(x^2-x+12) dx\]

OpenStudy (mathmale):

Suggestion: Factor x^2 - x + 12. Then apply a certain property of logs.

OpenStudy (danjs):

\[\large 9 \int\limits \ln(x^2-x+12)dx\] try letting u=ln(x^2-x+12) dv = dx

OpenStudy (danjs):

\[u = \ln(x^2-x+12)\] \[du = \frac{ 2x - 1 }{ x^2 - x +12 }dx\] dv = dx v = x

OpenStudy (mathmale):

Dan's suggestion is a good one. You could, alternatively, factor x^2-x+12, obtaining (x-4)(x+3). Note that the ln of (x-4)(x+3) is ln(x-4) + ln(x+3). You might (or might not) find this integrand easier to integrate. Try it both ways.

OpenStudy (danjs):

@mathmale That is a +12 on that polynomial.